### v1

# High-Precision Measurement of D(, ) Photodisintegration Reaction and Implications for Big-Bang Nucleosynthesis

Abstract
We report on a high-precision measurement of the D(, ) photodisintegration reaction at the newly commissioned Shanghai Laser Electron Gamma Source (SLEGS), employing a quasi-monochromatic -ray beam from Laser Compton Scattering. The cross sections were determined over =2.327–7.089 MeV, achieving up to a factor of 2.2 improvement in precision near the neutron separation threshold. Combined with previous data in a global Markov chain Monte Carlo (MCMC) analysis using dibaryon effective field theory, we obtained the unprecedentedly precise (, )D cross sections and thermonuclear rate, with a precision up to 3.8 times higher than previous evaluations. Implemented in a standard Big-Bang Nucleosynthesis (BBN) framework, this new rate decreases uncertainty of the key cosmological parameter of baryon density by up to 16% relative to the LUNA result. A residual 1.2 tension between constrained from primordial D/H observations and CMB measurements persists, highlighting the need for improved reaction rates and offering potential hints of new physics beyond the standard model of cosmology.
The hot Big-Bang theory was first proposed in 1946 by George Gamow [1], and is now the most widely accepted cosmological model of the universe. According to the Big-Bang theory, the universe began with a fireball approximately 13.8 billion years ago. Following inflation and cooling, primordial Big-Bang nucleosynthesis (BBN) began when the universe was approximately 3 minutes old (when the temperature was reduced down to approximately 1 GK), and ended less than half an hour later when nuclear reactions were quenched by the low temperature and density conditions in the expanding universe. Only the lightest nuclides were synthesized in appreciable quantities through BBN, approximately 75% 1H, 25% 4He, with a tiny amount of 2H, 3He and 7Li. These relics provide us with a unique window into the early universe [2, 3, 4].
In general, the primordial abundances of deuterium 2H (frequently written as D) and 4He inferred from observational data agree with the predictions of the BBN model, except for the cosmological lithium problem [3, 4]. Therefore, BBN has long been considered as one of the three historical pillars of the cosmological ‘Big-Bang’ model. Up to now, among all primordial light elements, deuterium is the most constraining one since both its astronomical observations and BBN predictions reach about a-percent precision. In particular, deuterium is a very fragile isotope that can only be destroyed after BBN throughout stellar evolution, and its most primitive abundance is determined from the observation of cosmological clouds at high redshift, on the line of sight of distant quasars. Most recently, the precision in deuterium observations in cosmological clouds has improved dramatically, reaching an accuracy of 1.19% for primordial deuterium abundance, i.e., D/H=(2.5270.030)10-5 [5]. In BBN, deuterium is produced through (, )D, and subsequently destructed by three nuclear reactions of D(, )3He, D(, )3He, and D(, )3H. These four reactions are the main sources of nuclear-physics uncertainty for the predictions of the primordial deuterium abundance [6, 7, 8, 9, 10]. Therefore, it is essential to determine their cross sections at the BBN energies with a similar high precision to further constrain the predicted deuterium abundance D/H as well as the key cosmological parameter of baryon density [11]. This makes it necessary to reevaluate robustness of the standard BBN model in the context of precision cosmology.
For the first primordial synthesizing reaction of (, )D (also referred to as [12]), Suzuki et al. [13] and Nagai et al. [14] measured the capture cross sections at =0.02, 0.04, 0.064 and 0.55 MeV with a precision of 5–10% by using the pulsed neutron beams; while for its inverse process D(, ) (also referred to as ), Bishop et al. [15] and Moreh et al. [16] measured the photodisintegration cross sections at 3 energy points of =2.504, 2.618 and 2.757 MeV by using the radioactive -ray sources with a precision of 4–7%; Hara et al. [17] measured at 7 energy points over =2.33–4.58 MeV with a precision of 6–10% by using the Laser Compton Scattering (LCS) -ray beams; other miscellaneous experimental information can be referred to Refs. [18, 19, 20, 21, 8, 22]. It should be noted that directly applying these experimental data to the BBN predictions would increase the uncertainties. The effective field theory (dEFT) calculations suggest that theory errors can be sufficiently smaller than the experimental ones. It can provide a very useful discriminant for theories and their perturbative schemes. The compilations of BBN reactions adopt these theory-based cross sections because they can provide more robust and accurate predictions than experiments alone. However, the uncertainties from the recent theoretical estimations of the cross section for at BBN energies are considerably different from each other, e.g., 4% in [23], 2–3% in [24], 1% in [25], and 1% in [12]. These differences could lead to different uncertainties in the BBN predictions. Therefore, more precise experimental data in the BBN energy region are strongly required to characterize the nuclear reaction model. In addition, predictions of the matrix theory at =0.1 and 1 MeV are found to deviate significantly from others by 4.6% [12], which also require experimental verification.
A new type of LCS -ray source, named Shanghai Laser Electron Gamma Source (SLEGS) [26, 27, 28], has been recently commissioned in China. SLEGS is one of the beamlines of Shanghai Radiation Synchrotron Facility (SSRF) [29]. It uses a CO2 laser [30] to interact with the 3.5 GeV electrons from the SSRF storage ring. This interaction can be implemented via both the slant-scattering and back-scattering modes, and generates quasi-monochromatic -ray beams in an energy range of 0.4–21.7 MeV, with a flux of 105–107 photons/s via a collimation system [31]. SLEGS is the first LCS facility to produce the high-flux -ray beam in a laser Compton slant-scattering mode (i.e., LCSS mode), which can provide more convenient energy scanning capability.
In this Letter, we report the results of a high-precision experiment for the D(, ) photodisintegration reaction at SLEGS. The photoneutron cross sections were measured at 22 energy points near the neutron threshold. Our new cross sections are up to a factor of 2.2 more precise than the previous ones [17]. The cross sections of (, )D have been evaluated by dEFT with a Markov chain Monte Carlo (MCMC) analysis, together with our new data and all other relevant experimental data. With the more precise cross sections evaluated, an unprecedentedly precise thermonuclear (, )D rate is thus obtained, approximately 1.9–3.8 times more precise than the previous ones [12, 8] in the BBN temperature regime. The impact of the present high-precision (, )D rate has been investigated with a standard BBN model, and astrophysical implications are discussed on the cosmological parameter of baryon density .
The experiment was carried out at the SLEGS facility, and the experimental setup is similar to that described in [32]. The quasi-monochromatic -ray beam was generated via the LCSS mode, achieved by the interaction of the electron beam from the SSRF and the CO2 laser. The laser was operated at a power of 5 W, with a frequency of 1 kHz and pulse width of 50 s [30]. The -ray beam was collimated by a 5 mm coarse collimator and a 2 mm fine one (C5T2) [31]. The collimated -rays irradiated a high purity D2O (99.9%) water target that was sealed in a pure aluminium container (=10 mm, =100 mm, with an equivalent deuterium areal density of =6.651023/cm2) at 22 different laser–electron collision angles. The weighted average -ray energies ranged from =2.327–7.089 MeV (in quasi-uniform energy steps) with a flux of approximately 105 photons/s. The residual -ray beam was attenuated by an external copper attenuator with a thickness ranging from 160 mm to 175 mm, which was monitored by a LaBr3 detector (34 inch) in real time. Similarly to Refs. [28, 32], the incident -beam energy spectra on the D2O target were reconstructed by those of the LaBr3 detector accordingly, as shown schematically in Fig. 1. Here, the target was aligned in the geometric center of a 4 flat-efficiency 3He neutron detector (FED) [33, 32], where the photoneutrons were moderated with the surrounding polyethylene material and subsequently captured by the 3He counters.
Similarly to the previous analysis method [34, 32], the measured photoneutron cross section for an incident -ray beam with maximum energy of can be expressed by:
Where is the normalized, =1, energy distribution of the -ray beam obtained using the method described in Ref. [28]. Here, represents the true monochromatic photoneutron cross section as a function of energy . is the number of photoneutrons detected by the FED, and is the number of photons incident on the target. is the average detector efficiency determined using the Ring-Ratio (RR) technique [33, 32]. is the number of target nuclei per unit area. = is a correction factor for a thick target, where is the mass attenuation coefficient, and is the target thickness. is the fraction of the -ray flux above the neutron separation energy of =2.2246 MeV [35]. Here, we define a weighted average -ray beam energy () as follows:
In fact, the quasi-monochromatic cross section introduced above is a folded or convoluted cross section over the whole energy range of the incident -ray beam, which is defined here as (also in Ref. [34]). The uncertainties in consist mainly of statistical, systematic and methodological ones. Thereinto, statistical uncertainties are within 0.5–1.6% for most data points, with only two data points closest to the neutron threshold having larger statistical uncertainties of 3.6% and 10%, respectively. The systematic uncertainty is mainly attributed to the following two factors [36, 32]: 1) the FED efficiency (3.0%); 2) the reconstructed incident -ray energy spectrum owing to the copper attenuator and the detection efficiency of LaBr3 (2.0%). The total systematic uncertainty is estimated to be approximately 3.6%. Furthermore, the methodological uncertainty [37] in the neutron extraction and spectral unfolding procedure is estimated to be 1.6–1.9%. For the total uncertainty, we have assumed that all the uncertainties estimated above are independent, and thus they were added quadratically. Table 1 lists the folded photoneutron cross sections and their uncertainties.
The cross sections available for the D(, ) photodisintegration can be well described by a phenomenological expression [22], i.e.,
with a parameterized Legendre polynomial coefficient
We have convoluted this parameterized cross section with the -beam energy spectrum and performed least-squares fits to the 22 experimental folded data points, and thus obtained the true monochromatic photoneutron cross sections () numerically (see details in End Matter). In this way, we obtain a continuous monochromatic D(, ) cross-section curve. Figure 2 compares our (unfolded) monochromatic photoneutron cross sections with the previous ones. It shows that our new values are a factor of 1.2–2.2 more precise than Hara et al.’s data [17], both with a similar experimental method. It should be noted that Hara et al.’s data are 5.2% systematically lower than the present ones (except for their lowest energy point), possibly because they are a kind of quasi-monochromatic or folded cross sections. Therefore, our precise measurements over a wide energy region (with better consistency) can act as a better reference for evaluating the (, )D cross sections. Table 1 summarizes the unfolded (true) monochromatic photoneutron cross sections and their uncertainties.
| Folded cross section (mb) | Unfolded cross section (mb) | |||||||
| 2.327 | 0.671 | 0.068 | 0.011 | 0.024 | 0.647 | 0.057 | ||
| 2.375 | 0.736 | 0.027 | 0.014 | 0.027 | 0.736 | 0.044 | ||
| 2.433 | 0.856 | 0.014 | 0.016 | 0.031 | 0.851 | 0.036 | ||
| 2.510 | 1.001 | 0.013 | 0.018 | 0.036 | 1.010 | 0.035 | ||
| 2.584 | 1.164 | 0.013 | 0.021 | 0.042 | 1.163 | 0.040 | ||
| 2.663 | 1.310 | 0.011 | 0.023 | 0.047 | 1.320 | 0.049 | ||
| 2.797 | 1.516 | 0.011 | 0.027 | 0.055 | 1.568 | 0.068 | ||
| 2.906 | 1.712 | 0.011 | 0.029 | 0.062 | 1.747 | 0.083 | ||
| 3.081 | 1.921 | 0.011 | 0.033 | 0.070 | 1.987 | 0.104 | ||
| 3.177 | 2.028 | 0.012 | 0.034 | 0.073 | 2.096 | 0.112 | ||
| 3.289 | 2.165 | 0.012 | 0.035 | 0.078 | 2.204 | 0.120 | ||
| 3.436 | 2.222 | 0.015 | 0.041 | 0.081 | 2.319 | 0.127 | ||
| 3.562 | 2.341 | 0.013 | 0.039 | 0.085 | 2.395 | 0.130 | ||
| 3.688 | 2.373 | 0.016 | 0.041 | 0.086 | 2.454 | 0.132 | ||
| 3.859 | 2.464 | 0.016 | 0.040 | 0.089 | 2.510 | 0.132 | ||
| 4.181 | 2.532 | 0.015 | 0.041 | 0.092 | 2.562 | 0.129 | ||
| 4.630 | 2.539 | 0.015 | 0.041 | 0.092 | 2.554 | 0.123 | ||
| 5.048 | 2.445 | 0.014 | 0.038 | 0.089 | 2.499 | 0.118 | ||
| 5.584 | 2.382 | 0.013 | 0.038 | 0.086 | 2.396 | 0.114 | ||
| 6.022 | 2.317 | 0.015 | 0.041 | 0.084 | 2.302 | 0.112 | ||
| 6.632 | 2.203 | 0.015 | 0.038 | 0.080 | 2.168 | 0.110 | ||
| 7.089 | 2.030 | 0.014 | 0.036 | 0.074 | 2.070 | 0.110 | ||
Similar to the method used in Ref. [12], a global fitting within the framework of the dEFT has been performed based on the Markov Chain Monte Carlo (MCMC) analysis. All the relevant experimental data, i.e., the scattering cross sections in the neutron energies of 10 MeV [39], the capture cross sections [13, 14], the photon analyzing power for the process [20, 21], and our folded cross sections of D(, ) obtained at SLEGS, have been included in the present calculations. The results of the global MCMC fitting are shown in Fig. 3. Notably, compared to the previous evaluation [12], both the present SLEGS data and recently measured scattering data [40] have been newly incorporated into this work, without including Hara et al.’s [17] data. Our evaluated capture cross sections are almost identical to the previous ones [12] (at most 0.5% difference), while the present uncertainties are significantly reduced to 0.11–0.28% in the energy region of 0.01–1.0 MeV, i.e., about 2.5–4.3 times more precise. Figure 4 shows the cross sections of (, )D (upper panel) and the corresponding ratios (lower panel), respectively. It shows that almost all experimental data are consistent with the present evaluation, but again Hara et al.’s data are 4.9% systematically lower for the reason discussed above. In addition, significant deviations can be observed for the -matrix results [24] (labeled with ‘Hale2001’) at 0.05, 0.1 and 1 MeV points (up to 4.6% at 0.1 MeV). The evaluation confirms that this -matrix calculation still needs certain modifications.
A new capture rate in the temperature range of 0.01–10 GK has been calculated with the presently evaluated (, )D cross sections, with the upper and lower limits of the rate estimated using the corresponding uncertainties of the cross sections. The uncertainty of the present rate is reduced significantly down to 0.12% in the temperature of BBN interest (see details in End Matter). In the BBN temperature region of 0.1–1 GK, the present rate (median value) is only slightly (at most 0.2%) larger than the previous Ando et al.’s evaluation [12], however, the corresponding uncertainty is reduced by a factor of 3.4–3.8; compared to the Serpico et al.’s rate [8], our rate is only slightly (at most 1%) smaller, but a factor of 1.9–3.2 times more precise. Thus, we recommend here an unprecedentedly precise rate for the (, )D reaction, which is of great importance for precision cosmological studies.
The impact of our high-precision (, )D rate has been investigated with a standard CDM BBN model [41, 42], with baryon density as a single free parameter, by performing a Bayesian analysis. Thereinto, the neutron life-time is taken as =879.40.6 s [43], and other 3 essential rates on determining D/H, namely, D, D and DHe, are taken from Ref. [44]. Using a deuterium abundance observed by Cooke et al. [5], i.e., D/H=(2.5270.030)10-5, we place a constraint of =0.022320.00033 in the standard CDM model with effective neutrino species =3.045 [45, 46], i.e., improving the uncertainty of by 10% compared to the previous LUNA value of 0.022330.00036 [47]. However, a tighter constraint of =0.022210.00031 can be achieved by a new observed value of D/H=(2.5470.025)10-5 from recent 11 quasar observations [48]. This will improve the uncertainty of by 16%.
It is known that the cross section of DHe measured by the recent LUNA experiment [47] has already reached an unprecedented precision of better than 3%. Thus, only DH and DHe reactions (hereafter short for and , respectively) now constitute the dominant nuclear physics uncertainties in predicting D/H and , because their cross section data still have large uncertainties. To evaluate the astrophysical -factors of and reactions for calculating their rates, Pisanti et al. [52, 44] employed polynomial fits to multiple experimental datasets, while Coc et al. [50] and Gómez et al. [49] utilized ab initio calculations as theoretical constraints for experimental data selection. The difference between the rates of Coc et al. and Gómez et al. is quite small (less than 0.4%), while those of Pisanti et al. deviate from other two rates up to factors of 6.3% and 3.3% for and reactions in the BBN temperatures, respectively. Thereinto, errors in the rates of Pisanti et al., Gómez et al. and Coc et al. were assumed to be 1.0%, 1.1% and 2.0%, respectively. Figure 5 shows the constraints on calculated using three different rates together with Cooke et al.’s observed value of D/H=2.5270.030)10-5 [5]. It shows that the values constrained with both Gómez et al. and Coc et al.’s rates only roughly match with the Planck CMB observations [51] (i.e., =0.022370.00015, indicated by a gray band). However, if a recently observed value of D/H=(2.5470.025)10-5 [48] is adopted, then a remarkable 1.2 tension in the constraints appears between the BBN prediction and the CMB observations once using the Gómez et al.’s rates. It shows that not only the rates (median value and associated error), but also the observed deuterium abundance can affect this tension.
In summary, we have studied the key BBN reaction (, )D via its time-reversal process D(, ) using a quasi-monochromatic -ray source. The cross sections obtained are much more precise than the previous ones, allowing evaluation of the (, )D cross sections and reaction rate with unprecedented high precision. Within a CDM BBN framework, this improvement reduces the uncertainty of the cosmological parameter by up to 16% compared to the previous result. The dominant nuclear physics uncertainties in the predictions of D/H and now arise from reactions, where different rate choices can lead to a 1.2 tension between constraints from primordial D/H observations and CMB measurements. Resolving such tension or discrepancy requires future high-precision measurements and theoretical refinements of reactions, which may reconcile current datasets or reveal new physics. This work also demonstrates the capability of SLEGS to deliver precise nuclear data of astrophysical importance.
Acknowledgments
We thank the staff of the Shanghai Synchrotron Radiation Facility (BL03SSID, https://cstr.cn/31124.02.SSRF.BL03SSID) for the assistance on this measurement. J. J. H. acknowledges Hiroaki Utsunomiya’s contribution to data reduction. This work was financially supported by the National Key R&D Program of China (Nos. 2022YFA1602401, 2022YFA1602400) and the National Natural Science Foundation of China (Nos. 12322509, 12275338, 12147101, 12335009, 12435010). Y. L. acknowledges the support of the Boya fellowship of Peking University and the China Postdoctoral Science Foundation (No. 2025T180924). S.-I. A. acknowledges the support of the National Research Foundation of Korea (NRF) grant (No. 2022R1F1A1070060 and 2023R1A2C1003177).